The sum of the measures of the angles in any triangle is 180 degrees. In triangle ABC, angle A is five more than twice the measure of angle B. Angle C is five less than three times the measure of angle B.
I tried: 2b+5=A 3B{c-5}
No luck

Answer by venugopalramana(3286)

The sum of the measures of the angles in any triangle is 180 degrees. In triangle ABC, angle A is five more than twice the measure of angle B. Angle C is five less than three times the measure of angle B.
I tried: 2b+5=A 3B{c-5}
GO STEP BY STEP AND GO SYSTEMATICALLY.
SINCE B HAS COMMON RELATION WITH BOTH A AND C..START WITH THAT
LET ANGLE B =B
LET US FIND A..
TWICE THIS =2B
5 MORE THAN THIS =2B+5=A
LET US FIND C...
THREE TIMES B..=3B
5 LESS THAN THIS ...
3B-5=C
SO NOW SUM UP THE 3 ANGLES...
A+B+C=2B+5+B+3B-5=180
6B=180
B=30
SO A=2B+5=2*30+5=65
C=3B-5=3*30-5=85